OpenStudy (anonymous):

When calculating a loan’s effective rate, if the interest compounds every two months, what value of n do you plug into your equation?

OpenStudy (anonymous):

n = number of times the interest is compounded per year

OpenStudy (anonymous):

a. 2 b. 0.167 c. 6 d. 60

OpenStudy (anonymous):

how many times would interest be compounded per year if every two months?

OpenStudy (anonymous):

im guessing 6 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank u Anna’s bank gives her a loan with a stated interest rate of 10.22%. How much greater will Anna’s effective interest rate be if the interest is compounded daily, rather than compounded monthly?

OpenStudy (anonymous):

a. 0.5389 percentage points b. 0.1373 percentage points c. 0.4926 percentage points d. 0.0463 percentage points

OpenStudy (anonymous):

D is answer

OpenStudy (anonymous):

My guess: effective interest rate: r=(1+i/n)^n -1 so for monthly r=(1+0.1022/12)^12 -1=0.107125762 and for daily r=(1+0.1022/365)^365 -1 = 0.107589126 To find the percentage points take compounded daily - compounded monthly: 0.107589126-0.107125762=0.000463364 to get percentage points: 0.000463364x100=0.0463 percentage points